OpenStudy (atianaaaa):

Evaluate the logarithm. log6 1/26

OpenStudy (photon336):

If I remember this correctly, we need to first re-write this. \[\log_{base} answer = exponent\] \[\log_{6} (1/26) => 6^{x} = (1/26)\] \[xlog(6) = \log(1/26)\] \[Log(1/26) \div \log(6) = x \]

OpenStudy (photon336):

@atianaaaa